OpenStudy (buttermequeasy):

6t-k=a t=? Undo the operations using SADMEP until you have isolated the variable. Remember, whatever is done to one side of the equation must be done to the other side.

OpenStudy (anonymous):

so add k to both sides then divide both sides by 6 6t -k + k = a +k 6t = a +k t = (a+k)/6